# ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM

### v1

@@ B.1 Baseline Models and Decoding Methods
LLaVA-1.5 7B is a minimalist yet efficient model that has served as the foundation for extensive studies in large multimodal modeling. LLaVA-NeXT 7B builds on LLaVA-1.5 with improvements in visual reasoning, higher input resolution, and enhanced world knowledge, resulting in superior performance on several benchmarks. Phi2-SigLIP leverages the Phi-2 backbone and a SigLIP-based vision tower, and is trained on the ShareGPT4V dataset, offering a compact alternative with competitive capabilities.
To assess the reliability and generalizability, we experiment with three decoding strategies: greedy search, nucleus sampling, and beam search. Greedy decoding yields deterministic outputs, while nucleus sampling and beam search enable for more diverse generation.
B.2 Datasets
With the widespread adoption of deep learning across computer vision and natural language processing Lu et al. (2023); Chen et al. (2025b, a); Cui et al. (2025); Yuan et al. (2025); Lu et al. (2024a, b); Zhao et al. (2025a, c); Xu et al. (2024, 2025), large language models (LLMs) have rapidly become foundational tools in AI systems Wang et al. (2024b); Ma et al. (2025); Wang et al. (2025b, a); Zeng et al. (2024); Xin et al. (2025); Du et al. (2021, 2025); Zhang et al. (2023c). However, despite their remarkable capabilities, MLLMs remain susceptible to hallucination—generating descriptions inconsistent with the input image. This phenomenon has sparked growing interest in evaluation benchmarks that specifically target hallucination in image-grounded generation.
@@ B.2 Datasets
B.3 Hyperparameters
The experiments are conducted with the following hyperparameter settings to obtain the best result:
$\kappa_{\textsc{vis}}=0.1$, $\alpha_{\textsc{neg}}=1.0$
LLaVA-1.5: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.6$, $\alpha=1.0$, and $\beta=0.5$.
LLaVA-NeXT: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.7$, $\alpha=1.0$, and $\beta=0.7$.
Phi2-SigLIP: $\kappa_{\textsc{tch}}=128$, $\alpha_{\textsc{pos}}=0.8$, $\alpha=0.5$, and $\beta=0.1$.
Appendix C Standard VQA Capability
Our approach substantially curbs hallucinations through explicit attention steering while simultaneously enhancing performance on standard VQA tasks, a trade-off that competing methods typically fail to avoid.
C.1 MM-VET
Method MM-VET Score
Orig 31.2
VCD 30.3
ICD 33.2
ASCD 33.2
Table 3: MM-VET scores for different methods.
C.2 ScienceQA
Method SQA Score
Orig 67.55
VCD 67.55
ICD 67.32
ASCD 69.51
Table 4: SQA scores for different methods.
C.3 GQA
Method GQA Score
Orig 61.28
VCD 59.38
ICD 59.99
ASCD 61.27
Table 5: GQA scores for different methods.
C.4 TextVQA
Method TextVQA Score
Orig 57.82
VCD 55.07
ICD 57.66
ASCD 57.91
Table 6: TextVQA scores for different methods.
C.5 MMMU
Method MMMU Score
Orig 0.342
VCD 0.333
ICD 0.360
ASCD 0.348
Table 7: MMMU scores for different methods.
Appendix D Visualization Example
Figure 7: Oringinal LLaVA-1.5 drifts completely, describing a riverside sidewalk, parked cars, and a stroller that do not exist in the image. VCD and ICD correctly identify the beach and ATV but still hallucinate extra objects such as a surfboard and a boat (highlighted in yellow). Our ASCD caption mentions only entities that are actually present and omits non-existent items, illustrating the improved visual grounding achieved through attention-steerable contrastive decoding.

### v2

@@ Appendix E Evaluation Settings
LLaVA-1.5 7B is a minimalist yet efficient model that has served as the foundation for extensive studies in large multimodal modeling. LLaVA-NeXT 7B builds on LLaVA-1.5 with improvements in visual reasoning, higher input resolution, and enhanced world knowledge, resulting in superior performance on several benchmarks. Phi2-SigLIP leverages the Phi-2 backbone and a SigLIP-based vision tower, and is trained on the ShareGPT4V dataset, offering a compact alternative with competitive capabilities. LLaVA-1.5 13B upsizes LLaVA-1.5 to 13B parameters—doubling language capacity while retaining the original architecture—so it serves as a clean test bed for model-scaling effects. Qwen2.5-VL-Instruct pairs the recent Qwen2.5-7B language model with a revamped vision encoder and multi-stage instruction tuning, representing a modern open-source architecture that achieves state-of-the-art performance out of the box.
To assess the reliability and generalizability, we experiment with three decoding strategies: greedy search, nucleus sampling, and beam search. Greedy decoding yields deterministic outputs, while nucleus sampling and beam search enable more diverse generation.
Datasets
With the widespread adoption of deep learning across computer vision and natural language processing, large language models (LLMs) have rapidly become foundational tools in AI systems. However, despite their remarkable capabilities, MLLMs remain susceptible to hallucination—generating descriptions inconsistent with the input image. This phenomenon has sparked growing interest in evaluation benchmarks that specifically target hallucination in image-grounded generation.
@@
Hyperparameters
The experiments are conducted with the following hyperparameter settings to obtain the best result on CHAIR:
$\kappa_{\textsc{vis}}=0.1$, $\alpha_{\textsc{neg}}=1.0$, $\alpha=1.0$, and $\beta=0.1$.
LLaVA-1.5: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.6$.
LLaVA-NeXT: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.7$.
Phi2-SigLIP: $\kappa_{\textsc{tch}}=128$, $\alpha_{\textsc{pos}}=0.8$.
Appendix F Visualization Example
Figure 8 compares the image–caption pairs produced by the original LLaVA-1.5, its VCD and ICD variants, and our ASCD (all with greedy decoding).
Figure 8: Original LLaVA-1.5 drifts completely, describing a riverside sidewalk, parked cars, and a stroller that do not exist in the image. VCD and ICD correctly identify the beach and ATV but still hallucinate extra objects such as a surfboard and a boat (highlighted in yellow). Our ASCD caption mentions only entities that are actually present and omits non-existent items, illustrating the improved visual grounding achieved through attention-steerable contrastive decoding.
$\alpha_{\text{pos}}$ Acc F1
Orig 85.4 84.1
0.1 86.0 85.1
0.2 85.9 85.1
0.3 85.8 85.0
0.4 86.0 85.3
0.5 86.3 85.9
0.6 86.6 86.3
0.7 85.2 84.2
$\alpha$ Acc F1
Orig 85.4 84.1
0.25 86.1 85.3
0.50 86.4 85.9
0.75 86.7 86.4
1.00 86.6 86.3
1.25 86.5 86.3
1.50 86.6 86.4
1.75 86.6 86.4
2.00 86.6 86.4
$\beta$ Acc F1
Orig 85.4 84.1
0.2 86.1 86.0
0.4 86.4 86.2
0.5 86.6 86.3
0.6 86.6 86.2
0.8 86.3 85.7
1.0 85.7 84.7
$\kappa_{\text{TCH}}$ Acc F1
Orig 85.4 84.1
16 86.3 85.7
32 86.6 86.3
64 86.6 86.3
128 86.5 86.2
256 86.3 85.7
512 86.0 85.4
Table 6: Ablation on key hyper-parameters (LLaVA-1.5 7B on POPE). Each block varies one hyper-parameter while fixing the others; bold marks the best setting per block.
Appendix G Runtime and Memory Overhead
Motivation.
Attention steering operates on internal matrices at generation time and could in principle introduce non-trivial latency or memory footprints. To verify that this is not a practical bottleneck, we benchmark ASCD against the two mainstream contrastive-decoding baselines—VCD and ICD—using the original LLaVA-1.5-7B model.
Setup.
For a single A100-80G we measure (i) wall-clock latency to produce an answer for a $224\times 224$ image–question pair and (ii) peak GPU memory, under three attention back-ends: the default PyTorch eager kernel, fused SDPA, and FlashAttention 2.
Metric Orig VCD ICD ASCD
Latency — eager (s) 0.09 0.17 0.17 0.16
Peak VRAM — eager (GB) 14.52 15.00 15.17 15.15
Latency — SDPA (s) 0.08 0.16 0.16
Peak VRAM — SDPA (GB) 14.52 15.15 15.17
Latency — FlashAttn (s) 0.08 0.15 0.15
Peak VRAM — FlashAttn (GB) 14.52 15.00 15.17
Table 7: Inference cost of different decoding strategies on LLaVA-1.5-7B. Dashes indicate that our current ASCD prototype can be integrated only in the eager path.
Observation.
In 7, ASCD adds at most $0.01$ s latency and $\sim$0.6 GB VRAM over the unmodified model, while VCD/ICD incur larger overhead due to image perturbations or prompt inflation. Thus, the selective head-level edits of ASCD preserve virtually the same inference efficiency, confirming that attention steering is a practical choice for real-time deployment.
Appendix H Ablation Study of Hyperparameters
Hyper-parameter robustness.
Table 6 sweeps the four controllable hyper-parameters of ASCD—positive-steer weight $\alpha_{\text{pos}}$, contrastive weight $\alpha$, truncation threshold $\beta$, and the number of text-centric heads $\kappa_{\text{TCH}}$—on the POPE benchmark with LLaVA-1.5-7B, whose mean accuracy and F1 jointly capture hallucination-reduction and VQA quality. Across the entire search range each variant consistently surpasses the untuned baseline, and the performance curve is flat around the optima (bold). The result indicates that ASCD is insensitive to moderate hyper-parameter changes and can be deployed with minimal tuning effort.